A balloon containing 0.143 mol Ne gas has a volume of 1.9 L at 1.9 atm and 28°C. How many grams of neon should be added to the balloon to increase the volume to 3.4 L at the same pressure and temperature?
Answer #1

let

x=moles of neon gas added

Total moles of neon now= (x+0.143)

Pressure = 1.9 atm and T= 28+273.15= 301.15K V= 3.4L

from PV= nRT

R=0.0821 L.atm/mole.K

n= PV/RT= 1.9*3.4/(0.0821*301.15) =0.2613

Total moles = x+0.143= 0.2613

x= 0.2613-0.143=0.1183 moles